Find the derivative of In(1+x^2) w.r.t tan inverse x
1
Expert's answer
2020-05-20T18:41:19-0400

Let us assume that:

"y = \ln (1 + x^2)"

And:

"v = \tan^{-1} (x)"

Derivative of "y" with respect to "x" is:

"\frac{dy}{dx} = \frac{d}{dx} \left( \ln (1 + x^2) \right) \\
\hspace{0.5 cm} = \frac{1}{1+x^2} \frac{d}{dx} (1 + x^2) \\"

"\hspace{0.5 cm} = \frac{1}{1 + x^2} \left[ \frac{d}{dx} (1) + \frac{d}{dx} (x^2) \right]"

"\hspace{0.5 cm} = \frac{1}{1+x^2} \left( 0 + 2x \right) \hspace{1 cm}" [ "\because \frac{d}{dx} (x^n) = nx^{n-1} \, and \, \frac{d}{dx} (c) = 0 \," ( Where c is a constant ) ]

"\hspace{0.5 cm} = \frac{2x}{1 + x^2}"

Derivative of "v" with respect to "x" is:

"\frac{dv}{dx} = \frac{d}{dx} \left( \tan^{-1} (x) \right) \\
\hspace{0.5 cm} = \frac{1}{1+x^2}" [ "\because \frac{d}{dx} \left( \tan^{-1} (x) \right) = \frac{1}{1+x^2}" ]

Now:

"\frac{dy}{dv} = \frac{\frac{dy}{dx}}{\frac{dv}{dx}} = \frac{\frac{2x}{1+x^2}}{\frac{1}{1+x^2}} = 2x"

Hence:

The derivative of "\ln (1 + x^2)" with respect to "\tan^{-1} (x)" or derivative of "y" with respect to "v" is:

"\frac{dy}{dv} = 2x"